OpenStudy (anonymous):

andy walked around the perimeter of the park 4 times the park is 500 feet by 700 feet. What is the total distance he walked

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

The perimeter of a plane figure is the distance around it.

OpenStudy (jadeishere):

9600 ft.... 500*2 = 1000 700*2 = 1400 2400 * 4 = 9600

OpenStudy (jadeishere):

I did that wrong didn't I?

OpenStudy (anonymous):

the answer choices: A.2500 B.5000 C.8000 D.10000

OpenStudy (jadeishere):

what the heck :/

OpenStudy (skullpatrol):

$$ \text{Perimeter of a rectangle = 2*length + 2*width}$$

OpenStudy (anonymous):

I know but it doesnt give a shape or figure

OpenStudy (skullpatrol):

"500 feet by 700 feet" means a rectangle

OpenStudy (skullpatrol):



OpenStudy (anonymous):

ok so can you give me some steps

OpenStudy (skullpatrol):

Calculate the distance around it.

OpenStudy (anonymous):

but if its a rectangle isnt all the sides going to be the same

OpenStudy (skullpatrol):

That would be a "square."

OpenStudy (skullpatrol):